A solution contains 2g of anhydrous bacl2 in 400cm3 it's conductivity is found to be 5.8*10-3 cm-1. Calculate the molar conductivity of bacl2 ( at. At.of bacl2 137)
 

Grade:12

3 Answers

Vikas TU
5 years ago
Molar conductivity is given as: K x 1000/M
M is the molarity.
M = moles of solute / Volume of solution in litres. => (2/137)/(400/1000) => 20/(137*4) => 0.0634
Molar conductivity => 5.8*10-3 * 1000/0.0634 => 5.8/0.0634 => 91.48 Sm^2/mol.
Boy Friend
2 years ago
Molar conductivity is given as :
∆m=conductivity*1000 / molarity
      =K*1000/M
Molarity= no.of moles of solute per liter of solution
And no. Of moles = mass in gram(x) / molecular wt.(W)
molecular weight of BaCl2 = 137 + (35.5*2) = 137+71=208
 
M= x / W*V(in litre)
    = (2/208) / (400/1000)                {400 cm3 = 400ml =                                                                         400/1000 Litre}
    =5/208
∆m= K*1000/M
      = 0.0058*1000 / (5/208)
      =5.8*208 / 5
      = 241.28 siemen cm^2 mol^-1
Boy Friend
2 years ago
There is some mistake in question hint 
Note that molecular mass of BaCl2 is 208 , not 137.
So i suggest every student to correct it and then solve your problem.
I hope this content is helpful for all of you.
                       Thanks
Yours 
Amit kr. Pandey